From: NASA HQ
Posted: Monday, November 28, 2005
SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2005) to enhance access to related status reports and NASA activities.
All ISS systems continue to function nominally, except those noted previously or below. Ahead: Week 8 for Expedition 12.
Following station inspection and morning hygiene, before breakfast and exercise, CDR/SO McArthur and FE Tokarev performed their third periodic (monthly) Russian biomedical assessment PZEh-MO-7 (calf volume measurement) and PZEh-MO-8 (body mass measurement), using the specially designed mass measurement device (IM), later breaking it down for stowage. [Calf measurements (left leg only) are taken with the IZOG device, a custom-sewn fabric cuff that fits over the calf, using the knee and lower foot as fixed reference points, to provide a rough index of deconditioning in zero-G and effectiveness of countermeasures. For determining body mass in zero-G, where things are weightless but not massless, the Russian IM "scales" measure the inertial forces that arise during the oscillatory motion of a mass driven by two helical metering springs with known spring constants. By measuring the time period of each oscillation of the unknown mass (the crewmember) and comparing it to the period of a known mass, the crewmember's mass is calculated by the computer and displayed.]
Today, Valery worked on the Russian STR thermal control system's KOB-1 loop to change out a pump assembly ORU (on-orbit replaceable unit, 3SPN1) against a new SPN from FGB stowage. [For the IFM (In-flight Maintenance), which was supported by a tag-up with ground specialists via VHF, the BITS2-12 onboard telemetry system and VD-SU control mode were powered down until the BITS-2-12 connector had been mated with the newly installed pump panel. A checkout of the assembly followed.]
CDR/SO McArthur had a number of science experiments on his work schedule, starting out with a familiarization review of procedures for today's and Wednesday's (11/30) checkout of the new Refrigerated Centrifuge (RC) in the HRF2 (Human Research Facility-2), delivered on LF-1/STS-114. [The subsequent checkout, which included installation of an RC retaining ring and setup, did not require power to the RC or HRF2 Rack.]
Afterwards Bill returned to the BCAT-3 (Binary Colloidal Alloy Test-3) experiment at the Lab MWA (Maintenance Work Area) to record the experiment hardware setup and to prepare an SSC (Station Support Computer) laptop for running the EarthKAM software for BCAT-3 operations prior to initializing Sample 6. [The equipment will take flash pictures automatically for the next three weeks, commanded from the EK camera control file. The behavior of these supercritical fluids is important because they combine the properties of liquids and gases. A better understanding of their reaction in the weightless environment of space could help in the development of new drugs, cleaner power, and interplanetary transportation.]
On the InSPACE (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions) coil assembly monitor experiment, McArthur inspected each assembly for possible damage, leaks and bubbles, as well as checked on the even dispersion of the particles in the magnetorheological fluid. [InSPACE obtains basic data on magnetorheological fluids, i.e., a new class of "smart materials" that can be used to improve or develop new brake systems, seat suspensions robotics, clutches, airplane landing gear, and vibration damper systems. The dispersed particles are contained in Coil Assemblies (CAs) which subject them to electric fields of certain strength and frequencies. Today's task was to monitor these particles.]
Valery Tokarev did the daily routine maintenance of the Service Module (SM)'s Environment Control & Life Support system (SOZh), including its toilet system (ASU), while Bill McArthur updated/edited the regular IMS delta file, including locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).
In the SM, the CDR performed the standard weekly maintenance on the TVIS (Treadmill with Vibration Isolation & Stabilization), which primarily checks the condition of the SPDs (Subject Positioning Devices) and records time & date values.
Both crewmembers completed their regular 2.5-hr. physical exercise program on the TVIS, RED resistive exerciser and VELO bike with bungee cord load trainer. [Valery's daily protocol prescribes a strict four-day microcycle exercise with 1.5 hr on the treadmill in unmotorized mode and one hour on VELO plus load trainer (today: Day 1 of the first set).]
Afterwards Bill transferred the TVIS and RED exercise data files to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM (Heart Rate Monitor) data from the workouts, followed by their erasure on the HRM storage medium (done six times a week).
At ~8:00am EST, CDR McArthur supported a PAO/TV event, downlinking a "Top 10" list for the "David Letterman Show" on CBS.
The Elektron oxygen generator continues to be off (since 11/26), to allow use of the remaining Progress O2 before 19P undocking on 12/20. [An 8.2 mmHg O2 repress was performed yesterday (11/27). Section 1 of the Progress tankage is now empty; ~69.5 kg of O2 remain in Section 2.]
19P propellant transfer to the ISS is scheduled for tomorrow.
TsUP/Moscow is looking at a possible delay in 19P undocking (from 12/20) to allow more time for consuming its stored O2 and conducting additional trash disposal. Ground teams are jointly studying this scenario, under which 20P would still be launched on time and would dock to the DC1 docking compartment's nadir port.
Two more GLA (General Luminaire Assembly) lighting fixtures have failed, one in the US Lab (leaving 10 functioning of 12), the other in the Node (leaving 7, of 8). [Power was removed from the GLAs by ground personnel per standard procedure, and repairs are being planned. There are seven spare LHAs (Lamp Housing Assemblies) and BBAs (Baseplate Ballast Assemblies) on board.]
No CEO (Crew Earth Observations) targets uplinked for today.
To date, over 177,000 of CEO images have been taken in the first five years of the ISS.
CEO photography can be viewed and studied at the websites:
See also the website "Space Station Challenge" at:
To view the latest photos taken by the expedition 12 crew visit:
Expedition 12 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/
Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.
ISS Orbit (as of this morning, 6:53am EST [= epoch]):
Upcoming Events (all dates Eastern):
Apogee height -- Mean Altitude -- Perigee height
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA's Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA's Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.
// end //